d) the element present in the greatest mass percent. e) the mass required to contain 1.00 g of P 6. A compound of carbon and hydrogen consists of 94.34% C and 5.66% H by mass. The molecular mass of the compound is found to be 178 amu. What is its molecular formula?
Answer #1

Element               %                 A.Wt     relative number            simple ratio                  simplest ratio

C 94.34 12 94.34/12 = 7.86        7.86/5.66   = 1.4          1.4*5 = 7

H 5.66               1                                   5.66/1      = 5.66          5.66/5.66   = 1               1*5   = 5

   Empirical formula = C7H5

molecular formula = (empirical formula)n

                     n        = M.Wt/E.F.Wt    = 178/89   = 2

molecular formula   = (C7H5)2 = C14H10